A coil of wire has a resistance of 33.8Ω at 28 °C and 44.6Ω at 69 °C. What is the temperature coefficient of resistivity?
Answer #1

The relation between resistance and coefficient of resistivity is given as

R = Ro (1 + (T - To))

R / Ro = 1 + (T - To)

just solve for , we get

= R / Ro - 1 /  T - To

= 44.6/33.8 - 1 / 69 - 28

= 7.793e-3 / oC